solve 7-2^x=-1. Add 24 to the value of x.
Answer by prateekagrawal(56)

7-2^x = -1
so,
8= 2^x
so,
2^3 = 2^x
x= 3.
add 24 to the value of x.
so, 24+3= 27 is the answer of this.